### v1

# Can a Quantum Computer Simulate Nuclear Magnetic Resonance Spectra Better than a Classical One?

Abstract
The simulation of the spectra measured in nuclear magnetic resonance (NMR) spectroscopy experiments is a computationally non-trivial problem, and as such, it represents a problem for which a quantum computer may provide some practical advantage over traditional computing methods. In order to understand the extent to which such problems may provide examples of useful quantum advantage, it is important to understand the limitations of existing classical simulation methods. In this work, we benchmark our classical solver designed to solve such problems. We find that it performs well, even beyond the common experimental parameter regimes, except for a specific molecule with certain unusual features. We discuss what implications this may have for future efforts to demonstrate quantum advantage in the context of NMR.
I Introduction
While advances continue to be made in the field of quantum computing hardware [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], demonstrating concrete examples of useful quantum advantage remains a challenge [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Considerable effort has been invested in searching for problems in which a solution provided by a quantum computer presents some advantage over a solution provided by a classical computer (including the extreme case in which a classical computer cannot reasonably provide a solution at all). One such problem which has been proposed as a candidate for demonstrating quantum advantage is the prediction of the spectra measured in nuclear magnetic resonance (NMR) spectroscopy experiments. Such a problem is seen as a natural candidate for quantum computation, as the Hamiltonian describing a molecule studied through NMR takes the form of a typical quantum spin system, the degrees of freedom of which map in a straightforward fashion to the qubits of a quantum computer. Previous work [34, 35, 36, 37] has investigated the manner in which such a problem can be solved through the use of a quantum computer.
Of course, understanding whether such a problem is truly capable of demonstrating useful quantum advantage requires understanding the limits of classical computation in the context of computing NMR spectra. In [38, 39, 40, 41, 42, 43, 44] the extent to which various classical approximation methods can tackle such a problem has been explored. In this work, we propose our own classical method for simulating NMR spectra, and examine its performance when applied to a selection of real-world molecules. We find that this solver provides accurate results in essentially all experimentally realistic parameter regimes, except for a small class of molecules with particularly unusual properties. We use the performance of our solver applied to these molecules to gain insight into the types of problems for which we may hope to see some degree of quantum advantage in the context of NMR spectroscopy.
The outline of this paper is as follows: In Section II we briefly discuss the basics of NMR spectroscopy, and in Section III we explain the details of our solver. In Section IV we discuss the performance of our solver when applied to a collection of molecules for which we possess an exact solution, and in Section IV we consider this performance for molecules for which we do not have an exact solution. We conclude in Section V. In Appendix A we explain how the particular spectral function we compute is related to quantities measured in an NMR experiment. In Appendix B we discuss how we compute certain accuracy metrics based on the output of our solver. In Appendix C we list where we obtain the parameters which define the molecules we study in this work. In Appendix D we display some additional supporting data which was omitted from the main text.
II Nuclear Magnetic Resonance
Nuclear magnetic resonance spectroscopy is an important tool in analytical chemistry, used to identify molecules and to elucidate their structure, among other applications [45, 46, 47, 48, 49, 50, 51]. In broad terms, NMR spectroscopy involves placing a sample of a given molecule in a large, constant magnetic field, perturbing it with a electromagnetic pulse, and measuring the response. In this section we provide a very brief discussion of the quantity which is measured in a typical NMR experiment, but in enough detail to elucidate the basic calculations which must be performed by our solver. We assume that the molecules are in a liquid solution, which simplifies some aspects of the analysis due to motional averaging.
The central object describing the physics of our molecule placed in an NMR spectrometer is the spin Hamiltonian (in radians per second) given according to
$\hat{H}=-\sum_{l}\gamma_{l}\left(1+\delta_{l}\right)B^{z}\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}$ (1)
The objects $\hat{I}^{\alpha}=\hat{S}^{\alpha}/\hbar$ with $\alpha\in\{x,y,z\}$ correspond to the usual SU(2) spin operators for a spin-S particle, which represent atomic nuclei present in the molecule. Only atomic nuclei with non-zero nuclear spin (“active nuclei”) will be relevant for an NMR experiment, with focus often being placed on nuclei with spin-1/2, such as 1H or 31P nuclei. The molecule is subjected to a magnetic field which we take to be in the Z direction, denoted by $B^{z}$, and each nuclei interacts with this field through a Zeeman interaction described by the gyromagnetic ratio $\gamma$. The different atomic nuclei can also couple to each other, parameterized by the terms $J_{kl}$. While each $J_{kl}$ could in principle be a tensor quantity, the motional averaging occurring in solution results in a single scalar quantity, thus resulting in an isotropic Heisenberg coupling. This term is generally a result of indirect spin-spin coupling mediated by electrons in chemical bonds.
The key fact which underpins the utility of NMR spectroscopy in general is that shielding of the external magnetic field by electrons in chemical bonds causes variations in the local magnetic field experienced by each nucleus in the molecule. In the case of molecules in solution, this can be characterized by the single scalar parameter $\delta$, the so-called chemical shift, which effectively modifies the gyromagnetic ratio of a given nucleus. The fact that different nuclei in a given molecule precess at slightly different Larmor frequencies,
$\omega_{l}=\gamma_{l}\left(1+\delta_{l}\right)B^{z},$ (2)
allows for the identification of different chemical groups present in the molecule. In proton NMR, which examines the NMR spectra of molecules that result primarily from the interaction of 1H nuclei with the external magnetic field, we define $\delta=0$ according to the effective gyromagnetic ratios experienced by the Hydrogen nuclei in the molecule tetramethylsilane (TMS), therefore setting the value of $\gamma$, as well as the values of the chemical shifts for the nuclei in any other given molecule (these gyromagnetic ratios in TMS happen to all be identical).
When a sample is placed in an NMR spectrometer, the magnetic field of the device results in some net magnetization of the sample. A brief electromagnetic pulse flips the spins into the XY plane, which therefore begin to precess, creating a signal that can be recorded by measuring the induced magnetic field along the X or Y axis. In Appendix A, we show that the measured X and Y components of this magnetic field directly lead to the so-called spectral function, which, up to proportionality, is defined as,
$C\left(\omega\right)\propto\eta\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (3)
where we have defined
$M^{\pm}\equiv\sum_{i}\gamma_{i}\left(\hat{I}^{x}_{i}\pm i\hat{I}^{y}_{i}\right)\equiv\sum_{i}\gamma_{i}\hat{I}^{\pm}_{i}\,.$ (4)
For the applicability of Eq. (3) see Appendix A. Note that we are typically not interested in the overall normalization of the spectral function, so in this work we will always normalize this function so that its integral is equal to the number of active nuclei in the molecule.
This (real and positive) quantity consists of a collection of Lorentzian peaks, centered at the transition frequencies of the system. The amplitudes of these peaks are determined by the matrix elements of the $M^{\pm}$ operators, which are restricted by the fact that the NMR Hamiltonian conserves the total spin along the Z axis,
$\hat{I}^{z}=\sum_{i}\hat{I}^{z}_{i}.$ (5)
This conservation allows the Hamiltonian to be written in a block-diagonal form, such that the largest block has dimension $\mathcal{D}^{2}$, with
$\mathcal{D}={\textstyle\binom{N}{N/2}}\;\sim\mathcal{O}\left((2^{N}/\sqrt{N})\right)\,$, (6)
where $N$ is the number of active nuclei in the molecule. If we were to attempt to calculate the spectral function directly by performing an exact diagonalization of the Hamiltonian on a classical computer, we would therefore anticipate that the required memory would scale as $\mathcal{O}\left(2^{2N}/N\right)$, while the required computational time would scale as $\mathcal{O}\left(2^{3N}/N^{3/2}\right)$. For specific molecules with highly symmetric coupling, one can additionally exploit a local SU(2) symmetry to reduce the effective system size [52]. For example, a freely rotating methyl group can be replaced by a decomposition into a spin-3/2 particle and two spin-1/2 particles, effectively decreasing the system size by one spin. However, the combinatorial effort associated with this reduction increases the complexity of the problem, and for small molecules (with less than around twelve nuclei), this can lead to an overall run time which is in fact longer. Furthermore, for molecules beyond a size of approximately twenty spins, neither of these symmetries is sufficient to obtain an exact solution in a reasonable amount of computational time, and further approximations must therefore be made. These approximations will be the basis for our solver, the details of which we discuss in the next section.
The quantity $\eta$, which we refer to as the “broadening” (sometimes known as the “line width”), is a convergence factor which is meant to model the rate at which the measured signal decays due to decoherence processes and to model a finite spectrometer resolution. In the non-interacting case, it is a straightforward exercise to verify that there is one peak corresponding to each nucleus, centered at the Larmor frequency of that nucleus. This quantity is therefore extremely useful in determining the chemical shifts of the molecule in question. In the presence of interactions, these peaks will typically split apart into additional smaller peaks, as the interaction terms delocalize the eigenstates of the system and allow for additional transitions, at slightly shifted frequencies. Therefore, it is generally desirable to operate NMR spectrometers at the largest possible magnetic field, since higher magnetic fields will enhance the effect of the Zeeman interaction compared with both the Heisenberg couplings and the broadening scale, leading to an NMR spectrum which can more easily identify the chemical shifts of the constituent nuclei. In this work, we will mostly focus on proton NMR. The parameter regimes in which typical NMR experiments operate result in a reference Larmor frequency $\omega_{\text{ref}}$ for the 1H nuclei in TMS which corresponds to a natural frequency $\nu_{\text{ref}}=\omega_{\text{ref}}/2\pi$ of a few hundred MHz, although smaller benchtop devices typically operate at lower frequencies, potentially reaching as high as 80 MHz. The chemical shifts in a typical molecule examined in proton NMR usually range from around $10^{-6}$ to $10^{-5}$ (often quoted as 1 to 10 “parts per million,” or “ppm” for short), resulting in peaks in the spectral function which are separated by anywhere from a few hundred to a few thousand Hz, while the coupling terms are typically on the order of tens of Hz. A broadening of a few Hz is usually reasonable. In our analysis in this paper, we will study three magnetic field strengths, corresponding to Larmor frequencies of 400, 80, and 20 MHz, which, for the purposes of this work, we will refer to as high, low, and very low field. We will also consider two different broadening values, corresponding to Lorentzian peaks with a full-width at half maximum of 1 Hz, and also 0.1 Hz, which we refer to as high and low broadening. Note that the parameter regimes corresponding to very low field or low broadening are somewhat atypical with respect to actual experiments and merely serve as a stress test for our solver.
III The Spin-Dependent Cluster Solver
To make a prediction for the spectral function which will be measured in an NMR experiment, we must compute the spectral function in equation (3). A brute-force solution to this problem would require either performing a time evolution of the Hamiltonian dynamics and then a transformation into the frequency domain, or an exact diagonalization of the Hamiltonian, which can be used to compute the sum in equation (3) directly. Due to the exponential scaling of the Hilbert space, such a brute-force approach will become very computationally expensive after about twenty spins, and computationally prohibitive after a few more, even if we do make use of the conservation of total spin along the Z axis and the SU(2) symmetry of symmetrically coupled groups.
To circumvent this issue, our solver is designed to compute the “spin-resolved” quantity
$C_{i}\left(\omega\right)=\gamma_{i}\sum_{j}\gamma_{j}\eta\sum_{n,m}\frac{\langle E^{(\Gamma_{i})}_{n}|\hat{I}^{-}_{i}|E^{(\Gamma_{i})}_{m}\rangle\langle E^{(\Gamma_{i})}_{m}|\hat{I}^{+}_{j}|E^{(\Gamma_{i})}_{n}\rangle}{\eta^{2}+[\omega-(E^{(\Gamma_{i})}_{n}-E^{(\Gamma_{i})}_{m})]^{2}}$ (7)
defined so that
$C\left(\omega\right)\approx\sum_{i}C_{i}\left(\omega\right).$ (8)
Dividing up the computation in this way, where we compute each $C_{i}$ separately, is advantageous because it allows us to take advantage of the key approximation utilized in the NMR solver, which is to construct a cluster $\Gamma_{i}$ of spins which are most likely to influence the physics of the chosen $C_{i}$, and then diagonalize the reduced Hamiltonian of this cluster, disregarding the effects from spins outside of the cluster. This approximation becomes exact in the limit where the cluster size coincides with the number of active nuclei. Note that these clusters can overlap, which is essential for obtaining accurate results. Furthermore, at fixed cluster size, this approach leads to a linear scaling with total molecular system size, though additional runtime improvements can further be made by diagonalizing the Hamiltonian for spins located in identical clusters only once, and also by exploiting the symmetry considerations outlined above, for each cluster Hamiltonian independently.
There are several heuristic justifications that one could provide for the potential success of such a clusterization method. In the time domain, if one imagines information (in the form of quantum correlations) spreading throughout the spin system after the pulse is applied, we expect that this information should only travel so far before the signal has decayed due to decoherence [53]. Thinking in the frequency domain, we expect that the effect of a small coupling term will be to slightly perturb the energies and eigenvectors of the Hamiltonian, leading to the locations and amplitudes of the peaks in the spectral function to be slightly altered, which may not be visible above the broadening scale. In any case, the ultimate validity of such an idea will be born out by its potential success in accurately computing these spectral functions.
To determine precisely which spins should be included in a cluster, we must rank them in order of importance. We expect the effect of a coupling $J_{ij}$ on the spectral function to be unimportant if $J_{ij}$ is small compared with the difference between the Larmor frequencies of those two spins. One metric for determining the relevance of a given spin when computing the quantity $C_{i}$ is thus given by the quantity
$\frac{J_{ij}^{2}}{\left|\gamma_{i}\left(1+\delta_{i}\right)B^{z}-\gamma_{j}\left(1+\delta_{j}\right)B^{z}\right|+\mathcal{E}}$ (9)
where $\mathcal{E}$ is a small term which prevents the metric from diverging in the case that two Larmor frequencies are identical. For the broadening values we study here, we find it sufficient to simply set $\mathcal{E}$ to a constant value of 0.1 radians per second, as altering this value does not make a noticeable difference in the resulting spectral functions. The user of the solver can, however, adjust this value if desired.
Figure 1: The convergence for selected molecules with 16 nuclear spins for the case of high broadening, for high, low, and very low field.
Figure 2: The convergence for selected molecules with 16 nuclear spins for the case of low broadening, for high, low, and very low field.
IV Exactly Solvable Examples
We begin by discussing the performance of our solver for a selection of molecules whose exact solution is accessible to us, both in terms of accuracy and computational resources. In this section and the next, the couplings and chemical shifts that we use come from a combination of experimental databases, the Spinach Library [43], and our own internal quantum chemical calculations (similar to the approach outlined in [54, 55, 56, 57]). In Appendix C, we provide a list of the molecules we study, and explain where we obtain the parameters for each of them.
To gauge the accuracy of our solver, we need a metric for describing the closeness of two functions. For this purpose, we will make use of the cosine similarity,
$\cos\theta_{ab}\equiv\frac{\int_{-\infty}^{+\infty}C_{a}\left(\omega\right)C_{b}\left(\omega\right)d\omega}{\sqrt{\int_{-\infty}^{+\infty}C^{2}_{a}\left(\omega\right)d\omega}~\sqrt{\int_{-\infty}^{+\infty}C^{2}_{b}\left(\omega\right)d\omega}},$ (10)
which essentially measures the “angle” between the two functions $C_{a}$ and $C_{b}$, as the integrals can be understood as inner products on the space of square-integrable functions. An important feature of this metric is that it does not change if one or both of the functions is scaled uniformly. This is a desirable property for our metric to have, since the essential information contained in an NMR spectrum does not change if it is scaled by a constant factor. In Appendix B, we discuss exactly how we compute this quantity from the numerical output of the solver. Since we will be mostly interested in small deviations between spectral functions, the specific error metric we will present is in fact the (base ten) logarithm of one minus the cosine similarity,
$\epsilon_{ab}\equiv\log\left(1-\cos\theta_{ab}\right).$ (11)
Figure 3: The convergence of the spectral function of Limonene as a function of cluster size, for low field and high broadening. The inverted horizontal axis is conventional in the chemistry literature.
To lend credence to the utility of this metric as a means of comparing spectral functions, and to give an intuitive sense for what different values of this metric mean, we display in Figure 3 the spectral function of the molecule Limonene at low broadening and low field, computed for several different maximum cluster sizes, in each case comparing against the exact result (which in this case corresponds to a molecule with sixteen nuclear spins). In these plots, rather than plotting against the angular frequency directly, we plot against the quantity
$\Delta=\frac{\omega-\omega_{\text{ref}}}{\omega_{\text{ref}}},$ (12)
where $\omega_{\text{ref}}$ is the Larmor frequency of protons. Note, in experiments the reference frequency is typically set by the resonance of the protons in tetramethylsilane. This shifted and scaled variable has peaks located at the values of the chemical shifts $\delta$, rather than the Larmor frequencies themselves, which makes the identification of chemical groups simpler, and is the convention in the NMR literature (this is sometimes referred to as “plotting against the chemical shift”).
As we increase the maximum cluster size from five to ten, and then again to fifteen, we see that the spectral function goes from relatively poor agreement to fairly good agreement, and then to almost perfect agreement (and in fact in the last case, we must zoom in significantly in order to see any deviation). These maximum cluster sizes correspond to error values of roughly -1.26, -2.62, and -5.24, respectively. This type of relationship between the numerical value of the error metric and the qualitative behaviour of the spectral functions, while admittedly somewhat subjective, is generally reflected across all of the molecules we have studied.
With this understanding of the error metric in mind, we present results for the convergence of three selected molecules as a function of maximum cluster size, all of which contain sixteen nuclear spins. We display the high broadening case in Figure 1 and the low broadening case in Figure 2. As we would expect, the convergence generally improves for higher field and higher broadening, with the broadening seemingly having the more pronounced effect on accuracy. It is true that the rate of convergence of the three molecules is not necessarily identical. However, we highlight the fact that there are no sudden jumps in accuracy, either up or down, as a function of maximum cluster size, and that the error of the approximation is (almost always) decreasing. This fact will be important for attempting to benchmark the performance of the solver when applied to molecules for which we do not have an exact solution.
In Appendix D we provide convergence data for additional selected molecules, which all display the same general behaviour.
Figure 4: The computational time (on a logarithmic scale) required for selected molecules with 16 nuclear spins, for the case of very low field and low broadening.
Figure 5: The memory (on a logarithmic scale) required for selected molecules with 16 nuclear spins, for the case of very low field and low broadening.
As for the required computational resources, in Figure 4 we display the required computational time to solve for the spectrum of the three molecules with sixteen nuclear spins, as a function of maximum cluster size, for the case of very low field and low broadening (other parameter regimes display similar behaviour). In Figure 5 we display a similar plot for the required memory. As anticipated, the required computational time and memory both grow exponentially, since we perform an exact diagonalization for each cluster. However, even the most computationally intensive calculation requires approximately one hour of computational time, and in many cases, exceptionally good accuracy is reached already at cluster sizes which only require seconds or minutes of computational time. For fixed cluster size, the scaling is linear with total system size, since each additional spin adds an additional cluster to be diagonalized (this time can be sub-linear if multiple spins correspond to the same cluster, reducing the number of diagonalizations which must be performed). We note that the varying run times and memory requirements for molecules of the same total size can be traced back to the number of magnetically equivalent spins in each molecule, as the solver automatically exploits the local SU(2) symmetry where applicable. This data was collected on an Intel Xeon W-2145 processor at 3.70 GHz using 8 threads.
Here we discuss the performance of the solver when applied to larger molecules for which we do not have an exact solution. Instead of comparing against an exact case, the solution for each cluster size is compared against the largest size we can simulate (referred to as the “reference” size). As we examine the spectral function corresponding to smaller cluster sizes, if we eventually reach a cluster size beyond which we no longer see any significant differences between the smaller size and the reference size, then we will consider the solutions to have converged. While this does not represent a rigorous proof of convergence, it is nevertheless plausible, given that we have not found any examples in the exactly solvable case in which the accuracy of the solution suddenly jumped, especially not after reaching a plateau.
Figure 6: The convergence for selected molecules with more than 16 nuclear spins for the case of high broadening, for high, low, and very low field.
Figure 7: The convergence for selected molecules with more than 16 nuclear spins for the case of low broadening, for high, low, and very low field.
In Figures 6 and 7 we display the results of our analysis for a selection of ten molecules of various size, all of them with a reference cluster size of seventeen. Once again, we see that the convergence is better for higher field and higher broadening, and the agreement between a cluster size of sixteen and a cluster size of seventeen is exceptionally good in all cases. The worst convergence is for the case of Friedelin at low broadening, with error values of -2.77, -2.70, and -2.75 for high, low, and very low field, respectively. We also notice that the error values for the molecule tert-Butyl(tert-butylphosphanyl)phosphane, also named 1,2-di-tert-butyl-diphosphane, (referred to here as “Diphosphane”) seem to progress in a somewhat erratic fashion, for several parameter regimes. We briefly investigate the spectral functions of both of these molecules.
IV.1 Friedelin
Figure 8: The molecule Friedelin, with fifty Hydrogen atoms.
Figure 9: The spectral function for Friedelin for the case of low field and low broadening.
Friedelin is a naturally occurring triterpenoid whose rigid friedelane skeleton contains five fused rings and a single carbonyl group at C-3, while the remaining thirty carbon atoms are fully saturated. Because every carbon atom (apart from the quaternary ring junctions) carries at least one hydrogen, the molecule hosts 50 distinct 1H nuclei. Simulating a molecule with 50 active nuclei pushes the boundary of what is possible with classical cluster approximations, and thus Friedelin serves as a useful “large molecule” benchmark for our NMR solver. In Figure 8 we show a schematic representation of the structure of the Friedelin molecule.
In Figure 9 we display the spectral function of Friedelin, for the case of low field and low broadening, which represents the worst convergence of any spectral function. While there are some small details which are not perfectly captured, the majority of the larger features appear to be resolved, to the extent that it is not clear whether a more precise simulation would add to the practical interpretation of the NMR spectrum. For the case of Friedelin, this implies that our solver is capable of reaching the level of accuracy necessary for industrial applications, especially since this level of broadening is already somewhat experimentally unrealistic. However, the lack of perfect agreement suggest that there may still exist some room for possible improvement from the use of a quantum device, if not for Friedelin, then for another, similarly large molecule.
IV.2 Diphosphane
Figure 10: The molecule Diphosphane, with two Phosphorous atoms and eighteen Hydrogen atoms.
Figure 11: The spectral function for Diphosphane, for the case of high field and low broadening.
The Diphosphane molecule is a single-bond analogue of hydrazine in which each phosphorus bears one tert-butyl group and one hydride (tBu–PH–PH–tBu). In Figure 11 we display the spectral function of Diphosphane at the two portions of the proton NMR spectrum where there is a non-zero signal. To understand the behaviour seen in the convergence plots, we examine the spectral functions for cluster sizes fourteen through seventeen (the largest cluster sizes in the convergence study), specifically for the case of high field and low broadening. We notice two large peaks at low chemical shift, which do not seem to change as a function of cluster size, and then a collection of narrow features at higher chemical shift. We note that the location of these features seems to oscillate as a function of cluster size, rather than converge (this behaviour in fact also occurs at lower cluster sizes).
To understand the origin of this behaviour, in Figure 10 we show a schematic representation of the structure of the Diphosphane molecule. This molecule consists of two central Phosphorous atoms each coupled to three Methyl groups, which together contain a total of nine Hydrogen atoms. Additionally, two Hydrogens are strongly coupled to the two 13P nuclei. As these couplings are direct and not mediated via the carbon atomes, the molecule features unusually high spin-spin interaction. Furthermore, the gyromagnetic ratio of the Phosphorous nucleus is roughly half that of the Hydrogen nucleus, implying that there are two very different energy scales at play in this system. Given these unusual features of the Diphosphane molecule, it is not surprising that a clustering approach may produce inaccurate results.
Figure 12: The spectral function for Diphosphane, for the case of high field and low broadening, at larger cluster sizes, which we can obtain due to the symmetry of the molecule.
To underline this conclusion, we can in fact compare to the exact solution for Diphospane, due to the high level of symmetry exhibited by this molecule. The two phosphorous atoms have the same chemical shift as each other, which is also true for the two central Hydrogen atoms. These two central Hydrogen atoms couple with the same strength to their respective phosphorous atoms. However, much more significant is the fact that all of the eighteen Hydrogen atoms contained in the Methyl groups have identical chemical shifts, and also couple identically to their respective phosphorous atoms. Thus, all of the “magnetically equivalent” spins can be combined into a single (higher representation) spin object, by exploiting the local SU(2) symmetry. Specifically, this reduction leads to an effective system of two spin-9/2 particles coupled to four spin-1/2 particles, which can be solved in approximately 10 minutes of computational time on a laptop with 32 Gigabytes of memory and an Intel(R) Core(TM) Ultra 7 165U processor.
In Figure 12 we again display the spectral function of Diphosphane at the two relevant portions of the spectrum, this time for larger cluster sizes. We continue to see the oscillating behaviour at higher chemical shifts, right up to the exact solution. In addition, we see a dramatic qualitative change in the spectrum at lower chemical shift as we increase the cluster size from 21 to 22 spins (the full size of the system). It is therefore indeed true that a clustering approach fails for such a molecule. However, it should be stressed that our method of solving for the exact spectrum of Diphosphane (which involves combining magnetically equivalent spins) only works well due to the high level of symmetry in this molecule, but for similar molecules with some level of disorder in their shifts and couplings, this specific approach would not be successful.
V Conclusion
In this work we have presented results for the performance of our NMR solver when applied to a collection of selected molecules. In almost every case, we have found good agreement for every parameter regime of practical relevance, with the performance increasing with higher magnetic fields and higher broadening. This demonstrates the utility of our clustering-based solver as a powerful tool for understanding molecular NMR spectra, even for molecules that we can not solve exactly.
The only case in which we see poor performance is the case of a phosphorous-containing molecule exhibiting unusually strong spin-spin interactions, due to the direct coupling of certain Hydrogen atoms to the 13P nuclei. While for this specific molecule we were able to calculate the exact spectrum due to its high level of symmetry, there may in fact exist molecules with similar (but not identical) structure, where this would no longer be possible. Such considerations point towards a possible first line of investigation in the direction of quantum advantage.
In general, the convergence of our solver becomes slower for decreasing broadening or lower magnetic fields. Is it then experimentally realistic to consider reaching even lower broadening values and magnetic fields? The answer to this question can be found in the field of zero field NMR spectroscopy [58], in which one makes use of polarized samples to measure the NMR spectrum at magnetic field strengths of only a few nanotesla, such that the Zeemann term can then essentially be neglected. Interestingly, the line width in these experiments can be on the order of 0.01 Hz. We therefore expect classical solvers (including our own) to struggle in this regime, leading to a second path towards potential quantum advantage [35].
We are actively investigating both of these approaches, and hope to report on them in more depth in the near future.
Acknowledgements.
This work was supported by the German Federal Ministry of Education and Research, through projects QSolid (13N16155) and Q-Exa (13N16065), as well as the European Innovation Council (EIC), through the Horizon Europe EIC Transition project HQS-NextNMR (grant agreement 101214526). We also want to thank Alexander Zech, Julius Kleine Büning, Peter Pinski, and Ilya Kuprov for helpful discussions.
Appendix A Derivation of the Spectral Function
Here we demonstrate how the spectral function
$C\left(\omega\right)\propto\eta\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (13)
is connected to measurements made in an NMR experiment, specifically the so-called pulse-acquire-experiment. To this end, we will first discuss what properties are measured in such an experiment, and then derive from them the spectral function.
A.1 The pulse acquire experiment
In an NMR experiment, the magnetic field resulting from the magnetization of the sample is measured. The magnetic moment of a single molecule in the sample along a specified direction is given by the expectation value of the appropriate spin operator,
$m^{\alpha}=\hbar\sum_{i}\gamma_{i}\langle\hat{I}^{\alpha}\rangle=\hbar\langle M^{\alpha}\rangle$ (14)
where we have ignored the small correction to the magnetic moment due to the chemical shifts, and have used the notation
$M^{\alpha}\equiv\sum_{i}\gamma_{i}\hat{I}^{\alpha},$ (15)
with $\alpha\in\{x,y,z\}$. As the induced magnetic field is proportional to the magnetic moment, it is sufficient to add up the contributions from the magnetic moments of all of the molecules in the sample, and multiply by the appropriate permeability to find the total magnetic field. If we assume that the sample is reasonably uniform, then the magnetic field will also be uniform, and will simply be some constant multiple of the magnetic moment of one molecule. Since we are not interested in the overall normalization of the spectral function, we can simply study this individual magnetic moment. Our task then is to understand how the quantity above evolves in time in an NMR experiment.
After a molecule is placed in an NMR spectrometer, it is allowed to come to thermal equilibrium, a state described by the density matrix
$\rho=\frac{1}{Z}e^{-\beta\hbar H}.$ (16)
with $Z$ the partition sum. For a spectrometer with a natural frequency $\nu$ of 500 MHz, the characteristic energy scale of the molecule is
$h\nu\approx 3.3\times 10^{-25}~\text{J},$ (17)
while the thermal energy scale is
$k_{B}T\approx 4.1\times 10^{-21}~\text{J}.$ (18)
Since the thermal energy scale is several orders of magnitude larger than the characteristic energy scale of the molecule, it is an incredibly good approximation to take
$\begin{split}Z\rho&=I-\beta\hbar H\\ &=I-\beta\hbar\left[-\sum_{l}\gamma_{l}B^{z}\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}\right],\end{split}$ (19)
where we have again ignored the correction due to the chemical shifts. Since the coupling terms are several orders of magnitude smaller than the Zeeman terms, the expectation values of the two-site operators are negligible, and we can simply take
$Z\rho=I+\beta\hbar B^{z}\sum_{l}\gamma_{l}\hat{I}^{z}_{l}=I+\beta\hbar B^{z}M^{z}$ (20)
The sample is then subjected to a sequence of magnetic pulses, which are engineered to induce a $\pi/2$ rotation along the X axis for each active nuclei present in the sample, the dynamics of which we take to be effectively instantaneous, so that the density matrix evolves according to
$\rho\to\exp\left(-i\frac{\pi}{2}\hat{I}^{x}\right)\rho\exp\left(+i\frac{\pi}{2}\hat{I}^{x}\right)$ (21)
where we have defined
$\hat{I}^{x}=\sum_{i}\hat{I}^{x}_{i}\,.$ (22)
and restricted the result to spin-1/2 systems. The extension to general spin-S systems is straightforward. Applying this transformation to the thermal density matrix, we have
$Z\rho_{0}\to I-\beta\hbar B^{z}M^{y}$ (23)
as the initial density matrix after the pulse.
After this pulse, the density matrix of the molecule will evolve in time according to
$\rho\left(t\right)=e^{-iHt}\rho_{0}e^{+iHt}\propto I-\beta\hbar B^{z}e^{-iHt}M^{y}e^{+iHt}.$ (24)
The magnetic moment is given according to
$m^{\alpha}=\hbar\langle M^{\alpha}\rangle=\hbar\text{Tr}\left[M^{\alpha}\rho\left(t\right)\right].$ (25)
Since $M^{\alpha}$ is traceless, only the second term in the density matrix contributes to the trace, and we have
$m^{x}=\hbar\langle M^{x}\rangle\propto\text{Tr}\left[M^{x}e^{-iHt}M^{y}e^{+iHt}\right],$ (26)
along with
$m^{y}=\hbar\langle M^{x}\rangle\propto\text{Tr}\left[M^{y}e^{-iHt}M^{y}e^{+iHt}\right],$ (27)
Furthermore, if we assume that this magnetization decays over time due to some internal decoherence processes, we should in fact send
$m^{\alpha}\to e^{-\eta t}m^{\alpha}\,.$ (28)
Note that in the derivation we have made use of a high temperature and high magnetic field expansion. However, due to the small energy scale of the NMR system, it is valid for temperatures above 1K and magnetic fields larger than the Earth’s magnetic field.
A.2 Computing the Spectral Function from the Correlation Function
We can interpret these formulas (26) and (27) for the magnetic moments as correlation functions, respectively denoted $C_{YY}$ and $C_{XY}$, making use of the notation
$C_{\alpha\beta}\left(t\right)=\text{Tr}\left[\left(M^{\alpha}\right)^{\dagger}e^{-iHt}M^{\beta}e^{+iHt}\right]$ (29)
If we extend this notation, and thereby define
$\begin{split}C_{++}\left(t\right)&=\text{Tr}\left[\left(M^{+}\right)^{\dagger}e^{-iHt}M^{+}e^{+iHt}\right]\\ &=\text{Tr}\left[M^{-}e^{-iHt}M^{+}e^{+iHt}\right],\end{split}$ (30)
then using the linearity of the trace, we can write
$C_{++}(t)=\left[C_{XX}(t)+C_{YY}(t)\right]+i\left[C_{XY}(t)-C_{YX}(t)\right].$ (31)
If we further make use of the rotational symmetry of the Hamiltonian under the transformation
$X\to Y~;~Y\to-X,$ (32)
we can rewrite this as
$C_{++}(t)=2\left(C_{YY}(t)+iC_{XY}(t)\right).$ (33)
Inserting two complete sets of states in the definition of $C_{++}$, we find
$\begin{split}C_{++}\left(t\right)&=\sum_{n,m}\langle E_{n}|M^{-}e^{-iHt}|E_{m}\rangle\langle E_{m}|M^{+}e^{+iHt}|E_{n}\rangle\\ &=\sum_{n,m}e^{-i\left(E_{m}-E_{n}\right)t}\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle\end{split}$ (34)
We now consider the half-sided transform of this object,
$C_{++}\left(w\right)=\int_{0}^{+\infty}C_{++}\left(t\right)e^{-i\omega t}e^{-\eta t}d\omega$ (35)
Note that the quantity $\eta$ serves here as a convergence factor, although its physical motivation has been discussed previously. Using the fact
$\int_{0}^{+\infty}e^{-i\left(\omega+E_{m}-E_{n}\right)t}e^{-\eta t}d\omega=\frac{-i}{\omega+E_{m}-E_{n}-i\eta}$ (36)
We find that
$C_{++}\left(w\right)=-i\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\omega+E_{m}-E_{n}-i\eta}.$ (37)
Taking the real part of this expression, we obtain
$\text{Re}\left[C_{++}\left(w\right)\right]=\eta\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\eta^{2}+[\omega+E_{m}-E_{n}]^{2}},$ (38)
which is the spectral function (13), as desired.
Appendix B Calculation of Cosine Similarity
Here we comment briefly on how we use the output of the solver to compute the cosine similarity.
While the user of our NMR solver can in principle specify the points at which to evaluate the spectral function, the default behaviour of the solver is to determine the most relevant frequency points automatically. This is done using an equal-area partitioning of the spectral function at different levels of approximation, resulting in a nonlinear frequency grid. This is advantageous numerically, since computational effort is not wasted on portions of the spectrum where there is no signal. By default, the solver will return the spectral function at 2,000 points, although for the atypically low line widths which we simulate in the low broadening regime, we find it necessary to increase this to 20,000 points. It is worth noting that this automatic identification of the relevant frequency points may result in different frequency values for the same molecule, depending on the specified parameters of the problem (magnetic field, maximum cluster size, etc.).
In order to compute the cosine similarity, we must compute the quantity
$\cos\theta_{ab}=\frac{\int_{-\infty}^{+\infty}C_{a}\left(\omega\right)C_{b}\left(\omega\right)d\omega}{\sqrt{\int_{-\infty}^{+\infty}C^{2}_{a}\left(\omega\right)d\omega}~\sqrt{\int_{-\infty}^{+\infty}C^{2}_{b}\left(\omega\right)d\omega}}.$ (39)
Each of the integrals in the above expression can be approximated with a Riemann sum as
$\int_{-\infty}^{+\infty}f\left(\omega\right)d\omega\approx\delta\sum_{i}f\left(\omega_{i}\right)$ (40)
where $\delta$ (not to be confused with the chemical shift) is the spacing between some equally-spaced points $\omega_{i}$ which cover the entire range in which at least one of the functions $C_{a}$ and $C_{b}$ is non-zero. This reduces the expression for the cosine similarity to
$\cos\theta_{ab}=\frac{\sum_{i}C_{a}\left(\omega_{i}\right)C_{b}\left(\omega_{i}\right)}{\sqrt{\sum_{i}C^{2}_{a}\left(\omega_{i}\right)}~\sqrt{\sum_{i}C^{2}_{b}\left(\omega_{i}\right)}}.$ (41)
For a sufficiently large number of points, this finite sum will converge to the value of the continuous integral.
In practice, we perform the integration with respect to the shifted and scaled variable,
$\Delta=\frac{\omega-\omega_{\text{ref}}}{\omega_{\text{ref}}},$ (42)
which does not alter the value of the cosine similarity, since the Jacobian of such a transformation is a simple constant factor, which cancels in the definition of the cosine similarity. We find that 100,000 frequency points is sufficient for the convergence of the cosine similarity in all of the cases we consider. Explicitly computing the value of the spectral function at all of these points is however unnecessary, as the values of the spectral function at the 20,000 points returned by the solver are sufficiently dense that a good interpolation of the data can be performed. In particular, we make use of the PCHIP shape-preserving interpolator in scipy, as this appears to be more robust against Runge phenomena at the edges of the spectrum, as opposed to other possible options. Any large gaps in the spectrum are also filled in explicitly with zero values, in order to aide the interpolation process.
Appendix C Parameter Source
In Table 1, we list the molecules which we study in our analysis, and indicate the source of the chemical shifts and couplings which appear in the NMR Hamiltonian. All of the active nuclei for the molecules mentioned below are 1H nuclei, except for the case of Diphosphane, which in addition to containing 20 1H nuclei, also possesses two 31P nuclei.
We divide the molecules into six groups, depending on how we obtain their chemical shifts and couplings. For the molecules in Group A, we compute the couplings internally, while we obtain the chemical shifts from the Biological Magnetic Resonance Data Bank [https://bmrb.io/]. For the molecules in Group B, we compute the couplings internally, while we obtain the chemical shifts from the publication [59]. For the molecules in Group C, the parameters are obtained from the collection of Hans Reich [https://organicchemistrydata.org]. For the molecules in Group D, we compute all of the parameters internally. We obtain the parameters for the Diphosphane molecule (Group E) from the Spinach library [43]. Lastly, for the molecule Friedelin (Group F), we compute the NMR parameters internally, although we have manually adjusted some of the shifts to match experimental data provided to us by PhytoLab GmbH & Co. KG.
Table 1: The list of molecules studied in this work, along with their size and parameter source.
Molecule Name Nuclei Count Molecule Group
Cyclobutane 8 B
THF 8 B
Cyclopentene 8 B
Cyclohexenone 8 B
Toluene 8 B
Norbornadiene 8 B
methylenecyclobutane 8 B
quadricyclane 8 B
Naphthalene 8 C
Propane 8 C
Cinnamaldehyde 8 D
t-Butyl chloride 9 B
2-Methyl-2-cyanopropane 9 B
Cyclopentane 10 B
t-Butylacetylene 10 B
MTBE (methyl t-butyl ether) 12 B
Anethole 12 D
Camphor 16 A
Limonene 16 D
Triphenylphosphine oxide 16 D
(+)-Neomenthol (SSR) 20 D
(-)-Menthol (RSR) 20 D
(+)-Isomenthol (SRR) 20 D
(-)-Neoisomenthol (SSS) 20 D
Artemisinin 22 A
1,2-di-tert-butyl-diphosphane (Diphosphane) 22 E
Androstenedione 26 D
Patchouli alcohol 26 D
4,10,11,11-Tetramethyltricyclo[5.3.1.01,5]undecan-10-ol 26 D
Friedelin 50 F
Appendix D Additional Convergence Data
Figure 13: The convergence for selected molecules with 8 nuclear spins for the case of high broadening, for high, low, and very low field.
Figure 14: The convergence for selected molecules with 8 nuclear spins for the case of low broadening, for high, low, and very low field.
Figure 15: The convergence for selected molecules with 10 nuclear spins for the case of high broadening, for high, low, and very low field.
Figure 16: The convergence for selected molecules with 10 nuclear spins for the case of low broadening, for high, low, and very low field.
Figure 17: The convergence for selected molecules with 12 nuclear spins for the case of high broadening, for high, low, and very low field.
Figure 18: The convergence for selected molecules with 12 nuclear spins for the case of low broadening, for high, low, and very low field.
In this Appendix we present some additional convergence data for exactly solvable systems. Figures 13 through 18 display the results for selected molecules of size 8, 10, and 12, for the usual parameter regimes. The convergence data is again generally what we would expect, although there are perhaps some stronger finite-size effects. We also have data for two 9-site molecules, t-Butyl chloride and 2-Methyl-2-cyanopropane, but both of these immediately converge to within machine precision for even the smallest cluster sizes, for every possible parameter regime.